# iRAG: An Incremental Retrieval Augmented Generation System for Videos

### v1

@@ IV Experimental results
We compare iRAG with the baseline video RAG system described in Section III. Baseline uses GRiT [10] for preprocessing, while iRAG uses DETR [7] and CLIP [12]. iRAG employs GRiT when preprocessed information cannot answer a user query. We compare the pre-processing time, average query processing time, and accuracy of the two systems. We briefly describe our video datasets in Table I, and the queries we pose on these datasets. We use outputs from the baseline system as the ground truth.
IV-A Datasets
@@
VQA-v2 validation dataset [13]: This dataset has 40,504 images. Of the 214,354 queries that are posed on these images, we selected 7799 queries, all of which start with “Is there”. We concatenate the images corresponding to these queries to create a long video (about 2 hours and 10 minutes long at 1 frame per second). We also randomly choose 1000 queries from the 7799 queries. These 1000 queries are used to evaluate iRAG and the baseline system.
@@
MSR-VTT [14]: This dataset has small clips of YouTube videos, with a caption per video. We concatenate several video clips to create a long video (about three and half hours long at 20fps). This dataset also has 318 queries.
@@
StreetAware: [11]: This dataset has 7.75 hours of video (2 TB of raw data) of traffic activity at four New York City traffic intersections. We selected a 46-minute video for our evaluation. This dataset does not have queries. As described later, we generated appropriate queries by suitably prompting an LLM.
@@
Tokyo MODI [15]: This dataset is a video of a busy street in Tokyo. Like others, we also used 2 hours of this video [15]. This dataset has no queries, and we prompt an LLM to generate queries for our evaluation.
TABLE I: Datasets used for evaluating iRAG
@@
IV-B Queries
Although two of our datasets (VQA-v2 and MSRVTT) come with queries, the real world datasets like StreetAware [11] and MODI [15] do not. Therefore, we use an LLM to generate appropriate queries. For this purpose, we ran the GRiT model [10] on both videos to generate captions, which are then used as context for the following prompt:
Prompt 2
@@ IV-B Queries
- Is there a white van?
- Is there a person with a backpack?
- Is there a bus?
@@
IV-C System implementation details
We implemented iRAG using the Langchain framework [16]. The vector database (with the FAISS index [17]) is from Langchain, and the LLM is OpenAI’s gpt-3.5-turbo model [18] API. We used DETR [7], an object detector, to do a fast preprocessing of videos for iRAG. In the incremental flow, we utilized GRiT [10], a popular image caption generator. However, this model requires significantly more computational resources (averaging approximately 1500 ms per frame) compared to the DETR object detector (averaging approximately 70 ms per frame). We conducted experiments on a server equipped with an AMD Ryzen 5950X 16-Core processor and an NVIDIA GeForce RTX 3090 GPU.
IV-D Evaluation metric
For the baseline and iRAG, we retrieve $k$ chunks as context ($k$ is a pre-determined configuration parameter for RAG). These chunks correspond to video clips. Given a query, we compute the ratio of the number of video clips common to both systems and the number of video clips in the baseline. If the ratio is 1, then both iRAG and the baseline retrieve the same video clips as context for the query. We define recall@k as the average of these ratios for all queries in a dataset. By varying $k$, we can compute different recall@k and study the impact of $k$ on the quality of context retrieved by iRAG (compared to the baseline).
@@ IV-D Evaluation metric
Table II shows the context quality recall@k for different values of $k$. For example, for the VQA dataset, when iRAG uses only DETR object detector for indexing (i.e. preprocessing), the recall@k for $k=20$ is $0.815$. This means that across all the queries in the dataset, there was on average about 81% overlap of video clips retrieved as context by iRAG and the baseline. Higher the value for recall@k, the closer are the responses from the LLM for the baseline and iRAG. However, naively increasing $k$ to improve context quality will also lead to an increased processing time in the incremental flow. Increasing $k$ can increase the number of video clips that will require detailed extraction during the incremental flow. This additional processing time can adversely affect interactive querying experience.
IV-E System evaluation
@@
We systematically evaluate the benefits from different enhancements in iRAG, compared to the baseline. We also study the impact of varying $k$ on the context quality and interactivity. This aids in the selection of a suitable $k$ for interactive querying.
@@
Benefit from using imageDB: By using lightweight vision-language models like CLIP, we can extract richer indexes in the preprocessing phase of iRAG. As shown in Table II, the context video clips retrieved by iRAG when both CLIP and DETR models are used in preprocessing have a better context quality than the case where only DETR is used.
@@
Performance of iRAG on the StreetAware and MODI datasets are shown in Figure 10 and Table VII, respectively. For these datasets, we show the human-understandable response video clips, and LLM responses from the iRAG system.
TABLE II: Performance of iRAG on VQA-v2 and MSRVTT datasets.
Dataset System Indexer r@1 r@2 r@4 r@8 r@20
@@
Benefit from chunk re-ranking: Results in Table II do not consider the case where the chunks retrieved by the Planner module are re-ranked by the Indexer to eliminate the less important chunks. The benefit from re-ranking context chunks is shown in Table III and Table IV. For the VQA-v2 dataset, with $k=8$, the Planner module retrieves 8 context chunks and the recall score is 66.07%. However, with chunk re-ranking, based on the query, the Indexer is able to eliminate less important chunks from the proposed 50 chunks from the Query Planner. Thus, the average number of chunks selected by the Indexer is only 6.88 while the recall score is 74.14%. Across all the 1000 queries for this data set, the average number of clips sent for detailed extraction in the incremental flow was only 6.88 chunks rather than 8 chunks. This results in reduced processing times for detailed extraction with accuracy improvement for the iRAG system with the refinement module. Therefore, iRAG suggests a cost-effective approach to extracting data for a query. For the classifier training, we select 1000 query from the training data of VQA-v2, for each query we run the GRiT model to extarct ground truth clips. Then we label the each concatenated query-chunk embedding vector as 1 if the chunk contains clips from ground truth, other wise give a 0. Then we train KNN classifier on this. After the KNN is trained, we test it on 1000 queries from the validation samples from VQA-v2 dataset to report results.
TABLE III: Benefit from re-ranking (VQA-v2 dataset)
@@
iRAG [w/o re-ranking] 8 8 0.6607
iRAG [w re-ranking] 50 6.88 0.7414
As the increase of k indicates the increase in total query processing time due to extracting unnecessary details using the Extractor, we define this k as a cost metric. We observe that for a specific k value (cost), the iRAG with re-ranking module offers improved accuracy (recall@k) score compared to iRAG w/o the re-ranking module. To visualize this, for iRAG, in Table IV, we set the initial k that the query planner offers are 10-50 chunk proposals. If we follow the query planner proposals, we may result in extracting frames from 10-50 chunks by the Extractor. Our iRAG with re-ranking module refines the k value and select it on average from $1.49\sim 6.88$ chunks. Now, if we observe the recall@k score for similar value of k ranging from 2-8 in iRAG w/o re-ranking, we observe that the iRAG with re-ranking improves the recall score by $1.9\%\sim 13.8\%$. Hence, the iRAG with re-ranking module is cost-effective and efficient.
TABLE IV: Impact of re-ranking on context quality
iRAG iRAG
w/o re-ranking with re-ranking
@@
(k) [after re-ranking] [before re-ranking]
@@
Preprocessing time:
Preprocessing of the long videos using DETR is a lot faster than preprocessing with GRiT. Table V shows that iRAG’s preprocessing time is a 23x to 25x less than the preprocessing time of the baseline.
TABLE V: Preprocessing time of iRAG and baseline
Data System Model Preprocessing Time
VQA-v2 iRAG DETR + CLIP 48 min 29 sec
RAG GRiT 1093 min 36 sec
MSRVTT iRAG DETR + CLIP 8 min 23 sec
RAG GRiT 199 min 50 sec
Query processing time vs $k$: Query processing time is important. It determines if the query-response session is interactive. It includes retrieval of the context clips by the Planner, optimizations in the Indexer, and detailed extraction of video clips by the Extractor. Figure 9 shows how interactivity (the query processing time) varies with $k$.
@@
Figure 9: Query processing time distribution for different values of $k$
Re-ranking context clips in the indexer eliminates less important clips, and improves interactivity. This is because due to re-ranking, not all the $k$ context clips are sent for detailed extraction. Re-ranking effectively reduces $k$ for the Extractor module, and the reduced detailed extraction processing time leads to significantly improved interactivity.
Note that the LLM query execution time remains almost constant across queries for both iRAG and the baseline. As a result, we have excluded this time from the query processing time calculations.
TABLE VI: Detailed extraction statistics
@@
2 0.437 0.323
4 0.578 0.392
6 0.670 0.428
8 0.724 0.445
10 0.771 0.458
20 0.907 0.493
30 0.968 0.503
40 0.986 0.508
50 0.995 0.509
Incremental extraction:
The baseline does detailed extraction of all the video clips during preprocessing. In contrast, iRAG does opportunistic, query-driven detailed extraction of clips as necessary. Table VI shows the fraction of the clips for which the Extractor performed detailed extraction for different values of $k$. For example, after processing all the queries in MSRVTT dataset, and even for large values of $k$, the Extractor performed detailed extraction for only about half the video clips.
@@
For VQA dataset, iRAG processes about 77% of the images for reasonable values of $k$ (between 2 and 10). The percentage is a higher because the long video was artifically created by us. Here, several images are packed into a video clip, and we do detailed extraction on clips rather than images. If we treat a video clip to be a frame, then the fraction of images processed by the Extractor is similar to the MSRVTT dataset case.
IV-F Human Evaluation
@@
Figure 10 shows frames in the context clips retrieved by iRAG and the baseline for a variety of queries. Visual inspection also shows that the context retrieved in iRAG is similar to the context retrieved by the baseline, except for one example. For the query, “Is there a white van?”, iRAG context is a poor match with the context from the baseline.
Table VII shows the LLM responses to queries in the Tokyo MODI dataset. In almost all cases, the responses from iRAG align well with the responses from the baseline system, except for the query “Are there any neon signs in Chinese in the video?”. The iRAG system does find clips with neon signs though. However, more measures are required to detect LLM hallucinations [19] and re-issue the user query.
Figure 10: Qualitative evaluation of iRAG system over StreetAware dataset.
TABLE VII: LLM API response on MODI street video dataset using iRAG
Query Incremental RAG (iRAG) RAG
Were there any black cars on the road? Based on the given context, there was a black car on the road at the time 1:52:10. Based on the given context, there was a black car on the road at 1:36:05. Therefore, the answer to the question “Were there any black cars on the road?” is yes, there was a black car on the road at 1:36:05.
Were there any people riding bicycles? Based on the given context, there is evidence of a person riding a bicycle in the following segment: ”When 1:37:15 I found the word shop on the building, people walking on the street, people crossing the street, the photo was taken in black and white, the people are walking, person walking across the street, white van parked on the side of the road, a large billboard, person wearing a jacket, person walking on sidewalk, a green building, person crossing the street, two people walking across the street, person on a bicycle, car parked on the street, person crossing the street.”Therefore, the answer to the question ”Were there any people riding bicycles?” is ”Yes, there was a person riding a bicycle.” Yes, there were people riding bicycles. In the context provided, it mentions ”person riding a bike” at both 1:58:15 and 1:12:50. Therefore, the answer to the question ”Were there any people riding bicycles?” is yes.
Were there any green trees lining the sidewalk? Based on the given context, I can confirm that there were green trees lining the sidewalk in the following instances: - When 1:52:10- When 0:59:30 - When 1:11:15 - When 1:44:30 - When 2:$0$ When 0:42:20 - When 0:29:45- When 0:10:35. Therefore, the answer to the question “Were there any green trees lining the sidewalk?” is “Yes, there were green trees lining the sidewalk in multiple instances.” Yes, there were green trees lining the sidewalk in both instances. The context provided mentions “green trees lining sidewalk” in both the timestamps 1:55:10 and 1:35:20. Therefore, the answer to the question ”Were there any green trees lining the sidewalk?” is yes.
Were there any neon signs in Chinese visible in the video? I can not answer. Yes, there were neon signs in Chinese visible in the video. The context provided mentions ”I found a neon sign in chinese” at both 0:28:50 and 0:33:35. Therefore, the answer to the question is ”Yes, there were neon signs in Chinese visible in the video.”
What color is the car on the street? Based on the given context, the color of the car on the street is black. This information is found in the context for the time stamp 1:52:10, where it states ”I found black car on the street.” Therefore, the color of the car on the street is black. The color of the car on the street is green and white. This information is found in the context provided, specifically in the segment ”When 1:28:25”. Therefore, the answer to the question ”What color is the car on the street?” is green and white.
Were there any people wearing hats? Based on the given context, I will search for instances of people wearing hats in the provided images. Let me analyze the images to find the answer to your question. Yes, there were people wearing hats in the given context. In the first timestamp (1:57:55), it mentions ”man wearing a black hat” and in the second timestamp (1:19:50), it mentions ”woman wearing a black hat.” Therefore, the answer to the question ”Were there any people wearing hats?” is yes.
Were there any green plants in the video? Based on the given context, the answer to the question ”Were there any green plants in the video?” is ”Yes, there were green plants in the video.” This is evident from the following excerpts from the context:- ”a green building” - ”a green tree” - ”a bunch of green plants” - ”a green tree in a parking lot” - ”green trees lining sidewalk”- ”a bunch of green produce” - ”a bunch of green flowers” - ”a planter with green leaves” - ”a green tree with lots of leaves” - ”a green tree in a planter” - ”a plant with purple flowers” - ”a planter with a lot of flowers” These excerpts provide clear evidence that there were green plants in the video. Yes, there were green plants in the video. At 1:03:55, green plants were found in a planter and green trees were lining the sidewalk. At 1:01:35, green plants were found in white vases.
@@
Were there any green trees with lots of leaves? Based on the given context, there are multiple instances of green trees with lots of leaves mentioned. Therefore, the answer to the question ”Were there any green trees with lots of leaves?” is ”Yes.” Yes, there were green trees with lots of leaves mentioned in the given context. The context specifically states ”I found a green tree with lots of leaves” in both instances at 0:19:35 and 0:15:35. Therefore, the answer to the question ”Were there any green trees with lots of leaves?” is yes.
Were there any umbrellas seen on the street? Based on the provided context, there is no mention of umbrellas being seen on the street. Therefore, the response to the question ”Were there any umbrellas seen on the street?” is ”No, there were no umbrellas seen on the street.” Based on the given context, there were umbrellas seen on the street. At 1:01:50, the umbrellas are described as green, and at 1:28:15, an umbrella is described as pink. Therefore, the answer to the question ”Were there any umbrellas seen on the street?” is yes, there were umbrellas seen on the street.
Were there any bicycles in the video? Based on the given context, I can not answer. Step 1: Analyze the given context for any mention of bicycles. Step 2: Review the context for any descriptions or sightings of a person riding a bike.Step 3: Identify any specific details related to bicycles, such as the presence of a bike or someone riding a bike. Step 4: Based on the analysis of the context, provide a response to the question. Response: Yes, there were bicycles in the video. At 0:05:40, a person was found riding a bike, and at 1:12:50, another person was found riding a bike.
Were there any green buses in the video? I can not answer. Yes, there were green buses in the video. At 0:35:30, the context mentions “green bus on the street” and “white writing on the bus.” Additionally, at 0:37:05, it states “the bus is green.” Therefore, based on the given context, we can confirm that there were green buses in the video.
@@
Were there any people wearing checkered shirts? I can not answer. Based on the given context, there were no mentions of people wearing checkered shirts. Therefore, the response to the question ”Were there any people wearing checkered shirts?” is “I can not answer”.
Were there any green traffic lights in the video? I can not answer. Based on the given context, there is no mention of any green traffic lights in the video. Therefore, the answer to the question “Were there any green traffic lights in the video?” is “No, there were no green traffic lights in the video.”
Were there any people wearing white jackets? Based on the given context, there is evidence of a person wearing a white jacket in the following segment: “When 1:54:30. I found red white and blue bus, a sign on a building, white van on street, person wearing a black hat, woman wearing a white jacket, black car on the street, the word shop on the building, the people in the city, a box of umbrellas, people walking on the street, person wearing a black jacket, green trees lining sidewalk.” Therefore, the answer to the question ”Were there any people wearing white jackets?” is “Yes.” Yes, there were people wearing white jackets. In the given context, it is mentioned that “woman wearing white jacket” was found at 1:45:35 and “person wearing white jacket” was found at 0:43:05. Therefore, the answer to the question “Were there any people wearing white jackets?” is yes.
V Prototype Implementation
Figure 11: Prototype implementation of the video iRAG system.

### v2

@@ 3. Experimental results
We compare iRAG with the baseline video RAG system described in Section 2. Baseline uses GRiT (Nguyen et al., 2022) for preprocessing, while iRAG uses DETR (Carion et al., 2020) and CLIP (Radford et al., 2021). iRAG employs GRiT when preprocessed information cannot answer a user query. We compare the pre-processing time, average query processing time, and accuracy of the two systems. We briefly describe our video datasets in Table 1, and the queries we pose on these datasets. We use outputs from the baseline system as the ground truth.
3.1. Datasets
@@
VQA-v2 validation dataset (vqa, 2024): This dataset has 40,504 images. Of the 214,354 queries that are posed on these images, we selected 7799 queries, all of which start with “Is there”. We concatenate the images corresponding to these queries to create a long video (about 2 hours and 10 minutes long at 1 frame per second). We also randomly choose 1000 queries from the 7799 queries. These 1000 queries are used to evaluate iRAG and the baseline system.
@@
MSR-VTT (msr, 2024): This dataset has small clips of YouTube videos, with a caption per video. We concatenate several video clips to create a long video (about three and half hours long at 20fps). This dataset also has 318 queries.
@@
StreetAware: (Piadyk et al., 2023): This dataset has 7.75 hours of video (2 TB of raw data) of traffic activity at four New York City traffic intersections. We selected a 46-minute video for our evaluation. This dataset does not have queries. As described later, we generated appropriate queries by suitably prompting an LLM.
@@
Tokyo MODI (Kossmann et al., 2023): This dataset is a video of a busy street in Tokyo. Like others, we also used 2 hours of this video (Kossmann et al., 2023). This dataset has no queries, and we prompt an LLM to generate queries for our evaluation.
Table 1. Datasets used for evaluating iRAG
@@
3.2. Queries
Although two of our datasets (VQA-v2 and MSRVTT) come with queries, the real world datasets like StreetAware (Piadyk et al., 2023) and MODI (Kossmann et al., 2023) do not. Therefore, we use an LLM to generate appropriate queries. For this purpose, we ran the GRiT model (Nguyen et al., 2022) on both videos to generate captions, which are then used as context for the following prompt:
Prompt 0.
@@ 3.2. Queries
3.3. Implementation details
We implemented iRAG using the Langchain framework (Harrison Chase, 2022). The vector database (with the FAISS index (Johnson et al., 2019)) is from Langchain, and the LLM is OpenAI’s gpt-3.5-turbo model (OpenAI, 2023) API. We conducted experiments on a server equipped with an AMD Ryzen 5950X 16-Core processor and an NVIDIA GeForce RTX 3090 GPU. For the classifier training, we select 1000 query from the training data of VQA-v2, for each query we run the GRiT model to extract ground truth clips. Then we label the each concatenated query-chunk embedding vector as 1 if the chunk contains clips from ground truth, other wise give a 0. Then we train KNN classifier on this. After the KNN is trained, we test it on 1000 queries from the validation samples from VQA-v2 dataset to report results.
3.4. Evaluation metric
@@
For the baseline and iRAG, we retrieve $k$ chunks as context that correspond to video clips. Given a query, we compute the ratio of the number of video clips common to both systems and the number of video clips in the baseline. If the ratio is 1, then both iRAG and the baseline retrieve the same video clips as context for the query. We define recall@k as the average of these ratios for all queries in a dataset. By varying $k$, we can compute different recall@k and study the impact of $k$ on the quality of context retrieved by iRAG (compared to the baseline). Higher the value for recall@k, the closer are the responses from the LLM for the baseline and iRAG. However, naively increasing $k$ to improve context quality will also lead to an increased processing time in the incremental flow. Increasing $k$ can increase the number of video clips that will require detailed extraction during the incremental flow. This additional processing time can adversely affect interactive querying experience.
3.5. System evaluation
@@
We systematically evaluate the benefits from different enhancements in iRAG i.e. preprocessing time, benefit of chunk filtering, query processing time, impact of varying $k$ on the context quality and interactivity compared to the baseline.
3.5.1. Preprocessing time:
Preprocessing of the long videos using DETR is a lot faster than preprocessing with GRiT. Table 2 shows that iRAG’s preprocessing time is a 23x to 25x less than the preprocessing time of the baseline.
Table 2. Preprocessing time of iRAG and baseline
Data System Model Preprocessing Time
VQA-v2 iRAG DETR + CLIP 48 min 29 sec
RAG GRiT 1093 min 36 sec
MSRVTT iRAG DETR + CLIP 8 min 23 sec
RAG GRiT 199 min 50 sec
3.5.2. Benefit from using ImageDB:
@@
By using lightweight vision-language models like CLIP, we can extract richer indexes in the preprocessing phase of iRAG. As shown in Table 3, the context video clips retrieved by iRAG when both CLIP and DETR models are used in preprocessing have a better context quality than the case where only DETR is used. Table 3 shows the context quality recall@k for different values of $k$. For example, for the VQA dataset, when iRAG uses only DETR object detector for indexing (i.e. preprocessing), the recall@k for $k=20$ is $0.815$. This means that across all the queries in the dataset, there was on average about 81% overlap of video clips retrieved as context by iRAG and the baseline.
@@
Performance of iRAG on the StreetAware and MODI datasets are shown in Figure 6 and Table 7, respectively. For these datasets, we show the human-understandable response video clips, and LLM responses from the iRAG system.
Table 3. Performance of iRAG on VQA-v2 and MSRVTT datasets.
Dataset System Models r@1 r@2 r@4 r@8 r@20
@@
3.5.3. Benefit from chunk filtering:
Results in Table 3 do not consider the case where the chunks retrieved by the Planner module are filtered to eliminate the less important chunks. The benefit from filtering context chunks is shown in Table 4 and Table 5. For the VQA-v2 dataset, with $k=8$, the Planner module retrieves 8 context chunks and the recall score is 66.07%. However, with chunk filtering, based on the query, the filtering classifier is able to eliminate less important chunks from the proposed 50 chunks from the Planner. Thus, the average number of chunks selected by the Planner after filtering is only 6.88 while the recall score is 74.14%. Across all the 1000 queries for this data set, the average number of clips sent for detailed extraction in the incremental flow was only 6.88 chunks rather than 8 chunks. This results in reduced processing times for detailed extraction with accuracy improvement for the iRAG system with the filtering module. Therefore, iRAG suggests a cost-effective approach to extracting data for a query.
Table 4. Benefit from filtering context clips (VQA-v2 dataset)
@@ 3.5.3. Benefit from chunk filtering:
iRAG [w/o filtering] 8 8 0.6607
iRAG [w filtering] 50 6.88 0.7414
As the increase of k indicates the increase in total query processing time due to extracting unnecessary details using the Extractor, we define this k as a cost metric. We observe that for a specific k value (cost), the iRAG with filtering offers improved accuracy (recall@k) score compared to iRAG w/o filtering. To visualize this, for iRAG, in Table 5, we set the initial k that the query planner offers are 10-50 chunk proposals. If we follow the context clips by Planner, we may result in extracting frames from 10-50 chunks by the Extractor. Our iRAG with filtering refines the k value and select it on average from $1.49\sim 6.88$ chunks. Now, if we observe the recall@k score for similar value of k ranging from 2-8 in iRAG w/o filtering, we observe that the iRAG with filtering improves the recall score by $1.9\%\sim 13.8\%$. Hence, the iRAG with filtering module is cost-effective and efficient.
Table 5. Impact of filtering on context quality
iRAG w/o filtering iRAG with filtering
@@
(k) [after filtering] [before filtering]
@@
3.5.4. Query processing time vs $k$:
Query processing time is important. It determines if the query-response session is interactive. It includes retrieval of the context clips by the Planner optimized by filtering, and detailed extraction of video clips by the Extractor. Figure 5 shows how interactivity (the query processing time) varies with $k$.
@@ 3.5.4. Query processing time vs $k$:
Figure 5. Query processing time distribution for different values of $k$
Filtering context clips by the Planner eliminates less important clips, and improves interactivity. This is because due to filtering, not all the $k$ context clips are sent for detailed extraction. Thus filtering context clips effectively reduces $k$ for the Extractor module, and the reduced detailed extraction processing time leads to significantly improved interactivity. Note that the LLM query execution time remains almost constant across queries for both iRAG and the baseline. As a result, we have excluded this time from the query processing time calculations.
3.5.5. Incremental extraction:
The baseline extracts detailed information of all the video clips during preprocessing. In contrast, iRAG does opportunistic, query-driven detailed extraction of clips as necessary. Table 6 shows the fraction of the clips for which the Extractor performed detailed extraction for different values of $k$. For example, after processing all the queries in MSRVTT dataset, and even for large values of $k$, the Extractor performed detailed extraction for only about half the video clips.
Table 6. Detailed extraction statistics
@@ 3.5.5. Incremental extraction:
k
Fraction of
video extracted
(VQA-v2)
Fraction of
video extracted
2 0.437 0.323 10 0.771 0.458
4 0.578 0.392 20 0.907 0.493
6 0.670 0.428 30 0.968 0.503
8 0.724 0.445 40 0.986 0.508
@@
For VQA dataset, iRAG processes about 77% of the images for reasonable values of $k$ (between 2 and 10). The percentage is a higher because the long video was artificially created by us. Here, several images are packed into a video clip, and we do detailed extraction on clips rather than images. If we treat a video clip to be a frame, then the fraction of images processed by the Extractor is similar to the MSRVTT dataset case.
3.5.6. Human Evaluation
Figure 6. iRAG’s qualitative evaluation on StreetAware videos (Piadyk et al., 2023)
@@ 3.5.6. Human Evaluation
Figure 6 shows frames in the context clips retrieved by iRAG and the baseline for a variety of queries. Visual inspection also shows that the context clips retrieved in iRAG is similar to the context clips retrieved by the baseline.
Table 7 shows the LLM responses to queries in the Tokyo MODI dataset. In almost all cases, the responses from iRAG align well with the responses from the baseline system. We also highlight a case in which the user query is ambiguous, specifically asking about the color of a car on the street. Both the baseline and our approach successfully retrieve a car, but they differ in the color identified.
Table 7. LLM response on MODI dataset (Kossmann et al., 2023)
Query Baseline iRAG Comments
Were there any black cars on the road? Yes Yes There was a black car on the road at the time at 1:52:10 and 1:36:05
Were there any people riding bicycles? Yes Yes There was a person riding a bicycle at 1:12:50, 1:37:15 and 1:58:15.
Were there any green trees lining the sidewalk? Yes Yes There were green trees lining the sidewalk at 0:10:35, 0:29:45, 0:42:20, 1:11:15, 1:35:20, 1:52:10, and 1:55:10.
Were there any people wearing hats? Yes Yes There were people wearing hats at 1:19:50, and 1:57:55.
@@
Were there any green trees with lots of leaves? Yes Yes There are multiple instances of green trees with lots of leaves mentioned at 0:19:35 and 0:15:35
@@
Were there any people wearing checkered shirts? I can not answer. I can not answer. Based on the given context, there were no mentions of people wearing checkered shirts
Were there any people wearing white jackets? Yes Yes There is evidence of a person wearing a white jacket in the following segment: 0:43:05, 1:45:35, and 1:54:30.
What color is the car on the street? Green and white Black In this case, both answers are correct as the time a black car is seen at 1:53:10 and green and white cars are seen at 1:28:15.
4. Prototype Implementation